1. A baby is 18 in. long at birth and 27 in. long at ten months. find the rate of change.

2. The cost of group museum tickets is $48 for four people and $78 for ten people. find the rate of change.

3. You drove 30 mi in one hour and 120 mi in four hours. find the rate of change.

4. (-7,1), (7,8) find the slope of the line passing through each pair of points.

5. (4,1 2/3), (-2, 2/3). find the slope of the line passing through each pair of points.

6. (0, 3.5), (-4, 2.5). find the slope of the line passing through each pair of points.

7. (1/2, 8), (1, -2). find the slope of the line passing through each pair of points.

8. (-5, 1/2), (-5, 3). find the slope of the line passing through each pair of points.

9. (0.5, 6.25), (3, -1.25). find the slope of the line passing through each pair of points.

10. Construction: An extension ladder has a label that says, "Do not place base of ladder less than 5 ft from the vertical surface." What is the greatest slope possible if the ladder can safely extend to reach a height of 12 ft? of 18 ft?

Answer by queenofit(120)

all of your questions can be solved using the following formula
y2-y1
_____
x2-x1

use points for all the problems. for problem 1 (0,18)and(10,27) are your points. Time always goes in the x spot. the first set of points are set x1 and y1 and the second set are x2 and y2. plug the numbers in the formula I gave and solve
27-18
-----
10-0
= 9/10
all are done the same way, Including slope. This is the formula for finding slope of a line and for finding rate of change when you have 2 points which you do in every question. I will do one more for you, number 3
(1,30)(4,120)
120-30
------
4-1
= 90/3 =30